How do I solve this one:
Sally was making a peanut butter and jelly sandwich. Sally cut off 1cm off each edge of the bread. If the bread is a square and the area removed was 32 cm^2 what is the length of a side of the sandwich (now that the 1 cm was removed)
I have tried using (x-1)(x-1) = 32 cm, but I already have been told an answer of "7" by my teacher. I cannot get the quadratic equation, when the coefficients are plugged in, to work and render '7'. Thanks.

Answer by stanbon(75887)

Let each side of the original sandwich have length "x".
When the 1cm is removed on all four sides the remaining
square has four sides that are x-2 cm long.
The area of the original sandwich is x^2.
the area of the remaining sandwich is (x-2)^2
EQUATION:
Original sandwich - New sandwich = 32 sq. cm.
x^2- (x-2)^2 = 32
x^2-x^2+4x-4=32
4x=36
x=9 cm (side of original sandwich before cutting)
x-2=7 cm (side of sandwich after cutting)
Cheers,
Stan H.